# From One Single Sketch to 3D Detailed Face Reconstruction

### v2

@@ Abstract.
copyright: noneccs: Computing methodologies Reconstructionccs: Computing methodologies Image processingccs: Applied computing Fine arts
Figure 1. A user creates a single facial sketch using our Sketch-1-to-3 system (left). The system produces a detailed 3D face reconstruction (right) that faithfully captures the sketch’s intended geometry and fine details, and is robust to occlusions, clutter, and diverse stylistic stroke variations.
[…]
The figure shows an interface of the Sketch-1-to-3 system on the left, where a user has drawn a simple free-hand facial sketch with black lines on a light blue background. To the right of the interface are three rendered views of the reconstructed 3D face model: a left-facing view, a frontal view, and a right-facing view. The reconstructed face is realistic, preserving the geometry and fine details implied by the sketch.
@@ 1. Introduction
We identify key limitations in sketch-based 3D face reconstruction, including the inherent ambiguity of hand-drawn sketches that hinders reliable information extraction, the difficulty of reconstructing high-fidelity 3D faces with vivid expressions and fine details from abstract line drawings, and the scarcity of high-quality facial sketch datasets for training. To overcome these challenges, we introduce Sketch-1-to-3, a novel framework that reconstructs high-fidelity 3D faces from a single sketch by accurately extracting contours and facial details. Furthermore, to address the data scarcity problem, we contribute two datasets: SketchFaces, a real hand-drawn sketch dataset encompassing diverse drawing styles, and Syn-SketchFaces, a synthetic dataset generated through a carefully designed synthesis pipeline.
Figure 2. Challenge demonstration. We employ(Lin et al., 2023) to transform the input sketch into the photo and then use(Feng et al., 2021) to generate the 3D face from the obtained photo, as elucidated by the yellow line. The green line signifies our method. It is apparent that employing the circuitous sketch-photo-3D (yellow line) method may give rise to reconstruction bias.
[…]
A diagram comparing two pipelines for 3D face reconstruction from sketches. The yellow line represents an indirect method that first converts the sketch into a photo and then reconstructs the 3D face from the photo. The green line represents our direct method. The comparison illustrates that the indirect method may introduce reconstruction bias.
@@ 3.1. Preliminaries
Figure 3. An overview of the proposed method, which consists of coarse and detail training stages. In the coarse stage, a shape encoder $E_{s}$ regresses $\bm{\beta}$ and a coarse encoder $E_{c}$ regresses $\bm{\theta}$, $\bm{\psi}$, $\bm{l}$, $\bm{c}$, and $\bm{\alpha}$. In the detail stage, a detail encoder $E_{d}$ generates a latent code $\bm{\delta}$ to refine the coarse 3D face with fine details. The 3D faces from both stages are rendered into 2D and compared with the input sketch to compute reconstruction losses. A GCTD module is applied in both stages to enhance feature extraction. All encoders employ inter-layer feature mixing for domain adaptation.
@@ 3.2. Enhancement Module of GCTD
The input sketch is first denoised using bilateral filtering, a non-linear, edge-preserving smoothing method that jointly considers spatial proximity $h_{s}$ and intensity similarity $h_{r}$. The process is defined as:
@@ 3.3. Loss Function
This Landmark loss calculates the correspondence of landmarks by compelling the projection of the 3D landmarks to align with the ground-truth landmarks of the sketch.
@@ 3.4. Domain Adaptation
However, the domain gap between synthetic and real sketches, which stems from differences in texture, stroke patterns, and abstraction levels, poses a significant challenge to the model’s generalization to real-world sketches. In addition, sketches from different sources or artists often exhibit substantial style variations, which further hinder generalization. Inspired by (Cai et al., 2024; Zhou et al., 2023), we incorporate a domain adaptation strategy into our approach, mixing feature statistics from the Syn-SketchFaces dataset and the SketchFaces dataset based on the inherent connection between image style and feature statistics(Huang and Belongie, 2017). During training, all encoders employ inter-layer feature mixing. We apply the mixing in lower layers, where features are most sensitive to texture, color, and stroke patterns. Let $F_{l}(\cdot)$ denote the encoder truncated at layer $l$. Given an input sketch $I_{\text{input}}$ and another sketch for mixing $I_{\text{mix}}$, we obtain the feature maps $x=F_{l}(I_{\text{input}})$ and $x^{\prime}=F_{l}(I_{\text{mix}})$. We then mix their channel-wise statistics at layer $l$ as
@@
This strategy not only bridges the gap between synthetic and real sketches but also enhances robustness to intra-domain sketch style variations, resulting in a marked performance gain on diverse sketch inputs.
Figure 4. (a) Syn-SketchFaces generation process: from original photos to PiDiNet contours to synthetic sketches. (b) Real sketch samples from the SketchFaces dataset.
[…]
Two subfigures showing datasets used in the study. Subfigure (a) presents the generation process of the synthetic dataset Syn-SketchFaces. It includes three columns: original face photos, contour maps extracted by PiDiNet, and the resulting synthetic sketches. Subfigure (b) shows samples from the SketchFaces dataset, which contains real hand-drawn facial sketches.
Figure 5. Qualitative comparisons. The left half shows reconstructions from synthetic sketches, and the right half from real hand-drawn sketches. Each group of five columns includes (from left to right): (1) input sketch, (2) DeepSketch2Face(Han et al., 2017), (3) SketchFaceNeRF(Lin et al., 2023), (4) the combination of(Lin et al., 2023) and(Feng et al., 2021), and (5) our Sketch-1-to-3.
[…]
A qualitative comparison of 3D face reconstruction results on real hand-drawn sketches.
@@
Table 3. User studies. Average ratings (scale: 1–5, poor to excellent) across four dimensions for expert (Group A) and non-expert (Group B) users.
@@ 4.3. Qualitative Comparisons
We conduct qualitative comparisons between our proposed method Sketch-1-to-3 and the current state-of-the-art 3D face reconstruction methods. The results demonstrate the superior performance of our method on both synthetic sketches and real hand-drawn sketches. See Figure 5. The five columns on the left illustrate the results of the 3D facial reconstruction in synthetic sketches, while the five columns on the right illustrate the results of the 3D facial reconstruction results on real hand-drawn sketches.
@@ 4.4. Ablation Studies
Figure 6. Ablation studies. In (a), the first row shows the input sketch, Sketch-1-to-3 without $L_{md}$, and the full model; the second row shows the input sketch, Sketch-1-to-3 without $L_{conP}$, and the full model. In (b), both rows show the input sketch, Sketch-1-to-3 without GCTD, and the full model.
[…]
Two subfigures showing results from ablation studies on the Sketch-1-to-3 model. Subfigure (a) includes two rows. The first row shows the input sketch, the reconstruction result without the $L_{md}$ loss, and the full Sketch-1-to-3 result. The second row shows the input sketch, the result without the $L_{conP}$ loss, and the full Sketch-1-to-3 result. Subfigure (b) shows two rows, each presenting the input sketch, the result without the GCTD module, and the full model output. These comparisons demonstrate the effectiveness of each component.
Ablation Studies on Loss Functions:
We conduct ablation studies by selectively removing the loss term during training. As illustrated in the first row, second column of Figure 6(a), excluding the mutual distance loss $L_{md}$ leads to poor alignment with the input sketch, particularly in the chin area. Additionally, in the second row, we investigate $L_{conP}$. Its removal causes deformation in the jawline, highlighting its role in preserving facial structure.
Ablation Studies on GCTD:
We conduct ablation studies by removing the GCTD module during training. In Figure 6(b), the second column shows reconstructions without GCTD, while the third column includes it. The first row demonstrates improved contour alignment and the second row reveals a more faithful detail reconstruction. Moreover, GCTD helps correct misalignments caused by recognition errors. These findings highlight the performance of the module in improving reconstruction accuracy.
@@ 4.5. User Studies
We conducted a user study to evaluate the real-world applicability of our method, comparing expert users with prior 3D modeling experience to non-expert users. Detailed design and procedures for the user study are provided in the Supplementary Material.

### v3

# Sketch-1-to-3: One Single Sketch to 3D Detailed Face Reconstruction

@@ Abstract.
journalyear: 2025copyright: acmlicensedconference: ACM Multimedia Asia; December 9–12, 2025; Kuala Lumpur, Malaysiabooktitle: ACM Multimedia Asia (MMAsia ’25), December 9–12, 2025, Kuala Lumpur, Malaysiadoi: 10.1145/3743093.3771049isbn: 979-8-4007-2005-5/2025/12submissionid: 393ccs: Computing methodologies Reconstructionccs: Computing methodologies Image processingccs: Applied computing Fine arts
Figure 1. A user draws a facial sketch (left), and our Sketch-1-to-3 system produces a detailed 3D face reconstruction (right) that faithfully preserves geometry and fine details, remaining robust to occlusions and stylistic stroke variations.:
@@ 1. Introduction
We identify key limitations in sketch-based 3D face reconstruction, including the inherent ambiguity of hand-drawn sketches that hinders reliable information extraction, the difficulty of reconstructing high-fidelity 3D faces with vivid expressions and details from abstract line drawings, and the scarcity of high-quality facial sketch datasets for training. To overcome these challenges, we introduce Sketch-1-to-3, a novel framework that reconstructs high-fidelity 3D faces from a single sketch by accurately extracting contours and facial details. Furthermore, we contribute two datasets, SketchFaces and Syn-SketchFaces, to address the data scarcity problem.
Figure 2. Challenge demonstration. The yellow line denotes a sketch-photo-3D pipeline using (Lin et al., 2023) and (Feng et al., 2021), while the green line represents our method. The indirect sketch-photo-3D process tends to introduce reconstruction bias.:
@@ 3.1. Preliminaries
Figure 3. An overview of the proposed method, which consists of coarse and detail training stages. In the coarse stage, a shape encoder $E_{s}$ regresses $\bm{\beta}$ and a coarse encoder $E_{c}$ regresses $\bm{\theta}$, $\bm{\psi}$, $\bm{l}$, $\bm{c}$, and $\bm{\alpha}$. In the detail stage, a detail encoder $E_{d}$ generates a latent code $\bm{\delta}$ to refine the coarse 3D face with fine details. The 3D faces from both stages are rendered into 2D and compared with the input sketch to compute reconstruction losses. A GCTD module is applied in both stages to enhance feature extraction. All encoders employ inter-layer feature mixing for domain adaptation.:
@@ 3.2. Enhancement Module of GCTD
The input sketch is denoised using bilateral filtering, an edge-preserving smoothing method considering spatial proximity $h_{s}$ and intensity similarity $h_{r}$. The process is defined as:
@@ 3.3. Loss Function
This Landmark loss calculates the correspondence of landmarks by compelling the projection of the 3D landmarks to align with the ground-truth landmarks.
@@ 3.4. Domain Adaptation
However, the domain gap between synthetic and real sketches, which stems from differences in texture, stroke patterns, and abstraction levels, poses a significant challenge to the model’s generalization to real-world sketches. In addition, sketches from different sources or artists often exhibit substantial style variations, which further hinder generalization. Inspired by (Cai et al., 2024; Zhou et al., 2023), we incorporate a domain adaptation strategy into our approach, mixing feature statistics from the Syn-SketchFaces dataset and the SketchFaces dataset. We apply the mixing in lower layers, where features are most sensitive to texture, color, and stroke patterns. Let $F_{l}(\cdot)$ denote the encoder truncated at layer $l$. Given an input sketch $I_{\text{input}}$ and another sketch for mixing $I_{\text{mix}}$, we obtain the feature maps $x=F_{l}(I_{\text{input}})$ and $x^{\prime}=F_{l}(I_{\text{mix}})$. We then mix their channel-wise statistics at layer $l$ as
@@
This strategy not only bridges the synthetic–real sketch gap but also enhances robustness to intra-domain sketch style variations, resulting in a marked performance gain on diverse sketch inputs.
Figure 4. (a) Syn-SketchFaces generation process: from original photos to PiDiNet contours to synthetic sketches. (b) Real sketch samples from the SketchFaces dataset.:
Figure 5. Qualitative comparisons. The left half shows reconstructions from synthetic sketches, and the right half from real hand-drawn sketches. Each group of five columns includes (from left to right): (1) input sketch, (2) DeepSketch2Face(Han et al., 2017), (3) SketchFaceNeRF(Lin et al., 2023), (4) the combination of(Lin et al., 2023) and(Feng et al., 2021), and (5) our Sketch-1-to-3.:
@@
Table 3. User studies. Average ratings (scale: 1–5, poor to excellent) across four aspects for expert (A) and non-expert (B) users.
@@ 4.3. Qualitative Comparisons
We conduct qualitative comparisons between our proposed method Sketch-1-to-3 and the current state-of-the-art 3D face reconstruction methods. See Figure 5. The five columns on the left illustrate the results of the 3D facial reconstruction in synthetic sketches, while the five columns on the right illustrate the results of the 3D facial reconstruction results on real hand-drawn sketches.
@@ 4.4. Ablation Studies
Figure 6. Ablation studies. In (a), the first row shows the input sketch, Sketch-1-to-3 without $L_{md}$, and the full model; the second row shows the input sketch, Sketch-1-to-3 without $L_{conP}$, and the full model. In (b), both rows show the input sketch, Sketch-1-to-3 without GCTD, and the full model.:
Loss Functions:
As shown in Figure 6(a), excluding $L_{md}$ leads to poor chin alignment, while removing $L_{conP}$ deforms the jawline, highlighting its role in preserving facial structure.
GCTD:
As shown in Figure 6(b), reconstructions with GCTD (third column) exhibit better contour alignment, finer details, and corrected misalignments, demonstrating its effectiveness in enhancing reconstruction accuracy.
@@ 4.5. User Studies
We conducted a user study to evaluate the real-world applicability of our method with expert users experienced in 3D modeling and non-expert users. Details are provided in the Supplementary Material.
@@ 5. Conclusion
Appendix A Overview
In this supplementary material, we provide additional details of our Sketch-1-to-3 framework, including implementation details, application scenarios, reconstruction results across diverse sketch styles, additional samples from the SketchFaces dataset, a user study description, and a discussion of limitations and future work.
Appendix B Implementation Details
During the first training stage, we conduct training for 10 epochs with a batch size of 16. In the second training stage, we set training for 15 epochs with a reduced batch size of 6. We employed Adam(Kingma and Ba, 2014) as the optimizer with a learning rate of 1e-4. We set $\omega_{eyeP}=10.0$, $\omega_{mouP}=5.0$, $\omega_{conP}=10.0$, $\omega_{pho}=0.2$, $\omega_{reg}=1e-05$. When calculating weighted landmark loss, we assigned weights to the inner mouth and the jawline as 3, the weight for the corners of the nose and mouth is set to 1.5, and the remaining parts are assigned a weight of 1. All three encoders share a ResNet-50 backbone(He et al., 2016), and we perform inter-layer feature mixing for domain adaptation.
Appendix C Application
The proposed method offers broad applicability in game development, animation, and digital human creation. By enabling sketch-based 3D face reconstruction, it allows artists to skip labor-intensive modeling steps and focus on creative design, while empowering hobbyists to produce high-quality 3D content with a lower entry barrier than traditional modeling tools.
Figure 7 illustrates the interface of our system, which provides an intuitive and flexible sketch-to-3D workflow. The interface supports multiple interactive controls. Users may draw a facial sketch directly within the panel or import an existing one created externally. With a single click, the system automatically generates a detailed and structurally coherent 3D face model from the input sketch, enabling both artists and non-experts to seamlessly transition from 2D concepts to high-quality 3D assets.
Figure 8 illustrates the application of our method in the 3D modeling pipeline. The reconstructed 3D face can be seamlessly imported into Blender for real-time preview and further editing, supporting a wide range of Blender’s built-in modeling and rendering tools.
Figure 7. Interface of our system.:
Figure 8. Application in 3D modeling tools.:
Figure 9. More results across various sketch styles. Our method robustly reconstructs 3D faces from sketches with diverse stroke styles, abstraction levels, and viewpoints, showing consistent geometry and stable performance across challenging cases. :
Figure 10. Limitation Demonstration. The model faces challenges when interpreting highly abstract or ambiguous sketch features, where complex facial details such as eyebrows or the mouth are reduced to a single dot or line. Such extreme simplifications can lead to uncertainty in geometric reconstruction and highlight the inherent ambiguity of sketch inputs. :
Appendix D Experiment
D.1. More Results Across Various Sketch Styles
As shown in Figure 9, we provide additional reconstruction results on sketches of diverse styles, further validating the robustness and generalization capacity of our Sketch-1-to-3 framework. The results show that our model accurately captures facial contours and fine details, while faithfully preserving head orientation, viewpoint, and expression, exhibiting stable performance under occlusions and diverse stroke styles.
D.2. User Study Design and Procedure
To assess the real-world applicability of our method, we conducted a user study involving 30 participants (17 males and 13 females) aged between 19 and 26. Participants are divided into two groups based on their 3D modeling experience: Group A consisted of individuals with prior training and practical experience in 3D modeling, while Group B had little to no such experience. Both groups had an equal number of participants.
Each participant was asked to independently draw five sketches of human faces. No specific stylistic restrictions were imposed, allowing participants to express themselves freely. The sketches were encouraged to include facial details, various head orientations, and occlusion to thoroughly test the abilities of our model. After processing the sketches through our model, each participant was shown the corresponding reconstructed 3D face results and asked to rate them on a scale from 1 (poor) to 5 (excellent).
Appendix E Limitation and Discussion
Although our method achieves promising results in general situations, there are still areas that can be explored in the future. The sketches inherently incorporate a pronounced level of ambiguity, posing a considerable challenge for the current model to accurately comprehend the intended information in sketches. As depicted in Fig 10, crucial features such as eyebrows and the mouth of the figure in the sketch can be succinctly represented by a single ink dot or a straight line, which we identify as a key factor contributing to the challenges in sketch-based 3D reconstruction.
Furthermore, as our work focuses on accurate facial geometry and detail reconstruction, we anticipate future efforts to deliver a complete 3D face creation pipeline. This would enable the reconstruction of colorized, texture-mapped 3D faces that faithfully reflect the artist’s intent. We think potential solutions to this challenge could come from colorization (Chen et al., 2024), generation (Zhan et al., 2023), and more.